OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality y 2x + 6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin

OpenStudy (aripotta):

y ? 2x + 6 which inequality is that?

OpenStudy (anonymous):

< with the line under it @AriPotta

OpenStudy (aripotta):

\[y \le 2x + 6 \]

OpenStudy (anonymous):

yes sir thats it @AriPotta

OpenStudy (aripotta):

ok, so we know it's going to be a solid line because it's less than or equal to

OpenStudy (aripotta):

the origin is (0,0) right?

OpenStudy (aripotta):

so just substitute the for the y and x in our inequality

OpenStudy (anonymous):

yessir @AriPotta

OpenStudy (aripotta):

substitute that*

OpenStudy (aripotta):

is 0 less than or equal to 2(0) + 6?

OpenStudy (anonymous):

idk :( im confused on this one

OpenStudy (aripotta):

what's 2 x 0?

OpenStudy (anonymous):

0

OpenStudy (aripotta):

we need to see if this is correct: \[0 \le 2(0) + 6\]

OpenStudy (aripotta):

\[0 \le 0 + 6\]

OpenStudy (aripotta):

is this true?

OpenStudy (aripotta):

is 0 less than/equal to 6?

OpenStudy (anonymous):

less than

OpenStudy (aripotta):

yea, so it's true.

OpenStudy (aripotta):

if it's true, then we know that the side that includes (0,0) is shaded

OpenStudy (aripotta):

so it does include the origin...

OpenStudy (anonymous):

so its a

OpenStudy (aripotta):

no...because earlier, we said that it would be a solid line, remember?

OpenStudy (anonymous):

oh right true so d

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

yeesss thank you i have more @AriPotta are you willing tohelp